# ABC: Achieving Better Control of Multimodal Embeddings using VLMs

### v1

@@ 4.5 VLM Backbone
In this section, we explore whether the choice of VLM backbone has an effect on the quality of our multimodal embedding. We ablate over $3$ popular choices of VLM backbone, all at approximately the $8$ billion parameter scale: Qwen2-VL-7B (Wang et al., 2024b), our chosen backbone. InternVL-2-8B from the internVL family of VLMs (Chen et al., 2024). Lastly, LLaVA-NeXT (Liu et al., 2024a) with Mistral-7B (Jiang et al., 2023) as its LLM component. We pretrain each model for $1000$ steps with a query batch size of $128$ and candidate batch size of $1024$.
Table 4: Accuracy vs. generative task performance.
@@
Table 4 shows the validation accuracy of ABC with different backbones. We find that our backbone choice, Qwen2-VL-7B, produces the best results. We also note each backbone’s performance on two standard generative VLM benchmarks: MMMU (Yue et al., 2024) and MMBench (Liu et al., 2024b). We find that performance after contrastive training strongly correlates with the performance of the backbone on generative tasks. This indicates that training better VLMs naturally results in better backbones for our embedding model.
5 Discussion and Future Work
@@ 5.1 Decoupling Pretraining and Instruction Fine-tuning
The separation of these two stages is an important step for increasing the accessibility of building instruction fine-tuned multimodal embedding models. Our pretraining run barely fits into the 640 GB of VRAM provided by a single A100 node, and takes several days to complete. Furthermore, our work indicates that these models could benefit substantially from further scaling (Appendix C). In contrast, the instruction fine-tuning stage completes in less than an hour. This allowed us to quickly iterate on our instruction fine-tuning stage, without pretraining from scratch. We hope that this decoupling encourages more practitioners to experiment with instruction fine-tuning methods for embeddings.
@@ 5.2 Important Factors during Pretraining
Prior work has largely focused on what architectural adaptations to the make the VLM to convert it into an embedding model. These include the attention mask, how the embedding is pooled and adapter architecture (BehnamGhader et al., 2024; Lee et al., 2024; Jiang et al., 2024b). Throughout our experiments, we find that most of these choices are often interchangable or only produce marginal improvements (Appendix B). However, We find that many of the crucial factors when training CLIP models are also important when adapting VLMs. In particular, well-chosen data, batch size and number of samples seen during training are all important factors (Appendix C), just like with CLIP models (Gadre et al., 2023; Cherti et al., 2023a).
5.3 Better Multimodal Benchmarks
Developing benchmarks that measure the capabilities of models to natural language instructions and image representations is an important direction for future work. We identify the following properties as crucial for a good quality benchmark. Firstly, the benchmark should require the use of both modalities together, and using only the image or the text should be insufficient to accomplish the task. We note that ensuring this property requires inspecting not only the queries but the candidate pool as well. For example, consider Figure 6, upon initial inspection the query seems well constructed. However, when the retrieval candidates are examined, it is clear that there is only one answer that is plausible given the instruction. Therefore, inspecting the image isn’t required to successfully complete the task. We find this a common pitfall when adapting open-ended generative benchmarks into multimodal embedding benchmarks.
Secondly, the instructions should be diverse. Users phrase natural language tasks in diverse and unpredictable ways, and benchmarks should reflect that (Trippas et al., 2024). It is easy to overfit on instruction phrasing, leading to misleading evaluations of the model’s ability to utilize instructions (Gudibande et al., 2023; Wei et al., 2022). Therefore, models should not be tested using instructions that they are trained on. Ensuring these two properties was our key motivation for constructing CtrlBench. However, CtrlBench only represents one multimodal embedding task, ideally more benchmarks satisfying the above two properties would be available to evaluate models.
Figure 6: An example of a poorly constructed task. The candidate pool has no relevant negatives, and therefore the image isn’t needed to solve the task.
@@ 6 Conclusion
We introduce ABC, a multimodal embedding model that leverages a VLM backbone to control image representations via natural language instructions. It achieves the best zero-shot results on a variety of multimodal tasks, spanning retrieval, classification, and VQA. Our multi-stage training process isolates the computationally expensive contrastive pretraining from a lightweight instruction finetuning phase, which allows for easy iteration of our model. We explore what factors are the most crucial when adapting VLMs to output multimodal embeddings. In particular, we find that training with well-chosen negatives, vision encoder resolution, and VLM backbone are all important factors for achieving the best performance. Lastly, we design CtrlBench to measure our model’s ability to use instructions to accomplish subtle natural language guided retrieval tasks.
Appendix A Instruction Prompt Design
@@
Appendix B Architecture Ablation
LLM2Vec (BehnamGhader et al., 2024) introduced several architecture modifications for adapting LLMs into text embedding models. In the multimodal setting, we find that using casual vs. bidirectional attention is marginal for improving model accuracy (Figure 7).
@@ Appendix B Architecture Ablation
On average, bidirectional attention slightly outperforms casual attention. Ablating over adapter rank, we find that different rank adapters tend to converge to the same accuracy, with higher rank adapters performing a fraction of a percentage better. Overall, better data and scaling (Appendix C) present a much more promising direction for improving embedding quality.
Appendix C Scaling Training
We find that batch size and number of samples seen during training are both important factors. By increasing our batch size by $4$x, from $128$ queries and $1024$ candidates to our full pretraining run size ($512$ queries and $4096$ candidates), our validation accuracy on an $800$ sample validation batch using in-batch negatives increases from $92.5\%$ to $95.0\%$. We control for total samples seen by training the smaller batch size model for $4$x more steps. Interestingly, this effect is even more pronounced on OOD validation data. For example, on a similar batch of MSCOCO data, accuracy increases from $73.8\%$ to $84.0\%$. Therefore, techniques for scaling batch size under limited VRAM, such as GradCache (Gao et al., 2021), are a promising direction to improve our pretrained model. Furthermore, we find that more steps (more samples seen) is also a straightforward way to increase the pretrained model’s performance. As shown by Appendix C, doubling step count steadily decreases loss during our pretraining run.
Appendix D Comparison to VLM2Vec
With the MMEB benchmark, Jiang et al. (2024b) also include a mutltimodal embedding model: VLM2Vec. We exclude VLM2Vec from Tables 1 and 2 as it has been trained on MSCOCO and several tasks from the MMEB benchmark. We provide the comparison of the two models here while specifying which tasks are out of distribution (OOD) for VLM2Vec. All tasks are OOD for our model.
@@ Appendix D Comparison to VLM2Vec
HatefulMemes (Kiela et al., 2021) 67.1 52.1
VOC2007 (Everingham et al., ) 88.6 81.4
SUN397 (Xiao et al., 2010) 72.7 71.8
Place365 (López-Cifuentes et al., 2020) 42.6 40.7
@@
ObjectNet (Barbu et al., 2019) 29.5 67.7
Country-211 (Radford et al., 2021) 13.0 18.5
@@
Table 5 compares the performance of our model to VLM2Vec (Jiang et al., 2024b). Gray rows indicate tasks that VLM2Vec is trained on. On average, VLM2Vec performs better on tasks that are in its distribution, while our model outperforms on tasks where both models are OOD. Interestingly, our model performs better on ImageNet-1K even though it is not trained on it.
Appendix E CtrlBench
  • Q: What is depicted in the image?
    A: This is a signage
  • Q: What can be seen growing near the base of the signpost?
    A: long limp green blades of plant
  • Q: What does the bottom sign say?
    A: a sign that says plums
  • Q: What is written on the top sign in this image?
    A: The word Apples written in white on a black background
  • Q: What does the middle sign in the image say?
    A: A dark grey sign that says Pears.
  • Q: What is featured prominently in the foreground of the image, facing the ocean?
    A: lone wood bench
  • Q: What do you notice about the water in this coastal scene?
    A: pretty blue ocean water
  • Q: What is the structure visible on the horizon in this coastal landscape?
    A: a distant, stone lighthouse
  • Q: What landmark is visible in the background by the ocean?
    A: old European lighthouse
  • Q: What do you notice about the weather in this seaside photo?
    A: blue sky in the distance

### v2

@@ 4.6 VLLM Backbone
We explore how the choice of VLLM backbone has an effect on the quality of our multimodal embedding. We ablate over $3$ popular choices of VLLM backbone, all at approximately the $8$ billion parameter scale: Qwen2-VL-7B (Wang et al., 2024b), our chosen backbone. InternVL-2-8B from the internVL family of VLMs (Chen et al., 2024). Lastly, LLaVA-NeXT (Liu et al., 2024a) with Mistral-7B (Jiang et al., 2023) as its LLM component. We pretrain each model for $1000$ steps with a query batch size of $128$ and candidate batch size of $1024$.
@@
Table 4: Ablation over VLLM backbone choice.
@@
Table˜4 shows the validation accuracy of ABC with different backbones. We find that our backbone choice, Qwen2-VL-7B, produces the best results. We also note each backbone’s performance on two standard generative VLLM benchmarks: MMMU Yue et al. (2024) and MMBench Liu et al. (2024b). We find that performance after contrastive training strongly correlates with the performance of the backbone on generative tasks. This indicates that training better VLLMs naturally results in better backbones for our model.
4.7 Controlling Embeddings using Natural Language
CLIP OpenCLIP SigLIP UniIR MagicLens BLIP2 MMRet ABC (ours)
VQA (10 tasks)
OK-VQA Marino et al. (2019) 7.5 11.5 2.4 25.4 12.7 8.7 28.0 48.1
A-OKVQA Schwenk et al. (2022) 3.8 3.3 1.5 8.8 2.9 3.2 11.6 37.3
DocVQA Mathew et al. (2021b) 4.0 5.3 4.2 6.2 3.0 2.6 12.6 28.5
InfographicsVQA Mathew et al. (2021a) 4.6 4.6 2.7 4.6 5.9 2.0 10.6 7.9
ChartQA Masry et al. (2022) 1.4 1.5 3.0 1.6 0.9 0.5 2.4 11.7
Visual7W Zhu et al. (2016) 4.0 2.6 1.2 14.5 2.5 1.3 9.0 25.6
ScienceQA Lu et al. (2022) 9.4 10.2 7.9 12.8 5.2 6.8 23.3 26.3
VizWiz Gurari et al. (2018) 8.2 6.6 2.3 24.3 1.7 4.0 25.9 29.4
GQA Hudson & Manning (2019) 41.3 52.5 57.5 48.8 43.5 9.7 41.3 60.1
TextVQASingh et al. (2019) 7.0 10.9 1.0 15.1 4.6 3.3 18.9 35.4
All VQA 9.1 10.9 8.4 16.2 8.3 4.2 18.4 31.0
Table 5: Zero-shot VQA results on MMEB (Jiang et al., 2024). Best is bold, second best is underlined.
Table˜5 shows ABC’s performance on the VQA split of MMEB. In just 100 training steps using instructions, our model surpasses MMRet (Zhou et al., 2024) on VQA, which was trained with millions of instructions. Interestingly, CLIP derivatives fine-tuned with instructions such as MagicLens (Zhang et al., 2024) and UniIR (Wei et al., 2023) do not perform significantly better than their respective baselines, despite MMEB providing instructions for each task. This evidences that encoder-only embedding models struggle to effectively utilize natural language instructions (Jiang et al., 2024).
Problems with VQA. A VQA benchmark should require the use of both modalities together, and using only the image or the text should be insufficient to accomplish the task. We note that ensuring this property requires inspecting not only the queries but the candidate pool as well. Consider Figure˜6, an example from the Visual7W task in MMEB. When the candidate pool is examined, it is clear that there is only one answer that is plausible, given the image. Therefore, the instruction is not required to accomplish the VQA task. This a common pitfall when adapting open-ended generative VQA benchmarks into fixed candidate pool embedding tasks.
Figure 6: An example of a VQA task from the Visual7W (Zhu et al., 2016) split of MMEB (Jiang et al., 2024).
CtrlBench. We construct CtrlBench, an instruction-controlled retrieval benchmark. CtrlBench has a similar format to the Flicker30K (Plummer et al., 2016) test split. However, instead of each image having multiple valid captions, we instead provide an instruction from which the model can infer the most relevant text candidate. Therefore, CtrlBench tests both retrieval and instruction following capabilities. To construct CtrlBench, we sample a $1000$ images from ADE20K (Zhou et al., 2017). To create $5000$ instructions and text candidate pairs, we generate $5$ instructions for each image, each corresponding to a distinct aspect of the image. To accomplish this, we use Gemini 2.5 Flash (Team et al., 2025). We provide the exact prompt that we use in Appendix˜B. As each image has $5$ associated captions in the candidate pool, a model that cannot utilize instructions can (in expectation) achieve at most $20\%$ R@1 on the benchmark. Therefore, CtrlBench requires the embedding model to use the provided instructions to disambiguate visual queries. We deduplicate captions from the dataset to prevent ambiguity in our text candidates, resulting in CtrlBench being slightly smaller than 1000 images.
Model R@1 R@5 R@10
UniIR Wei et al. (2023) 0.9 2.59 4.30
MagicLens Zhang et al. (2024) 17.0 37.33 48.47
VLM2Vec Jiang et al. (2024) 2.95 8.49 12.36
GME-Qwen2VL-2B Zhang et al. (2025) 20.92 46.33 57.27
GME-Qwen2VL-7B Zhang et al. (2025) 24.75 54.48 65.27
ABC (pretrained) 11.71 32.48 42.69
ABC (finetuned) 50.94 78.43 87.47
Table 6: Recall@K on CtrlBench.
Table˜6 shows the performance of multimodal embedding models on CtrlBench. We compare with $2$ instruction fine-tuned CLIP derivatives: UniIR Wei et al. (2023) and MagicLens Zhang et al. (2024) as well as VLM2Vec Jiang et al. (2024) and GME (Zhang et al., 2025), two concurrent works that adapt VLLMs into multimodal embedding models. We find that neither of the CLIP-based architectures have above $20\%$ R@1 on CtrlBench, the performance that indicates that the model is non-trivially utilizing instructions. We find that VLLM architectures are better at instruction-controlled retrieval. ABC and VLM2Vec have R@1 above $20\%$, proving that they can non-trivially utilize instructions.
5 Discussion and Related Work
5.1 Distribution Overlap in Training and Evaluation
Throughout our experiments, we found that several training data sources overlap with commonly used benchmarks. Several benchmark are constructed from the same underlying data source as training data or directly derived from training data (Barbu et al., 2019; Russakovsky et al., 2015; Plummer et al., 2016; Lin et al., 2015). This makes evaluating “zero-shot” capabilities difficult. To the author’s knowledge, our pretraining data (Sharma et al., 2018) and instruction-finetuning data Krishna et al. (2016) does not overlap with any benchmarks we evaluate against, and all evaluations (Tables˜1, 2, 5 and 6) are out-of-distribution.
@@ 5.2 Decoupling Pretraining and Instruction Fine-tuning
We find that adapting an instruction fine-tuned VLLM to use instructions is relatively cheap. Our instruction fine-tuning completes in less than an hour on a couple A100 GPUs. This allowed us to quickly iterate on our instruction fine-tuning stage, without pretraining from scratch. Comparatively, training a VLLM to create high-quality embeddings is much more resource intensive. Our pretraining run barely fits into the 640 GB of VRAM provided by a single A100 node, and takes several days to complete. Furthermore, our work indicates that these models could benefit substantially from further scaling (Appendix˜D).
5.3 Alternative Approaches
MagicLens (Zhang et al., 2024) and UniIR (Wei et al., 2023) are multimodal embedding models that fuse CLIP embeddings. By combining multiple CLIP models, they take image-text pairs and align them with image-text pairs using contrastive loss and instruction finetuning. Due to late fusion of modalities, these architectures struggle to use instructions to modify representations (Jiang et al., 2024). Recently, VLM2Vec Jiang et al. (2024) and GME (Zhang et al., 2025) have also adopted contrastive training on VLLM backbones to produce multimodal embeddings. However, these approaches make extensive use of human-annotated instruction datasets, with hundreds of thousands of labeled examples used during training. This limits the scalability and adaptability of their method.
@@ 5.4 Important Factors during Pretraining
Prior work has largely focused on what architectural adaptations to the make the VLLM to convert it into an embedding model. These include the attention mask, how the embedding is pooled and adapter architecture (BehnamGhader et al., 2024; Lee et al., 2024; Jiang et al., 2024). Throughout our experiments, we find that most of these choices are often interchangeable or only produce marginal improvements (Appendix˜C). However, we find that many of the crucial factors when training CLIP models are also important when adapting VLMs. In particular, well-chosen data, batch size and number of samples seen during training are all important factors (Appendix˜D), just like with CLIP models (Gadre et al., 2023; Cherti et al., 2023a).
@@ 6 Conclusion
We introduce ABC, a multimodal embedding model that leverages a VLLM backbone to control image representations via natural language instructions. It achieves strong zero-shot results on a variety of multimodal tasks, spanning retrieval, classification, and VQA. ABC consists of a multi-stage training process, which isolates the computationally expensive contrastive pretraining from a lightweight instruction finetuning phase. We explore what factors are the most crucial when adapting VLMs to output multimodal embeddings. In particular, we find that vision encoder resolution, contrastive loss temperature and VLLM backbone choice are all important factors. Lastly, we design CtrlBench to measure our model’s ability to use instructions to accomplish subtle natural language-guided retrieval tasks.
7 Broader Impact Statement
The ability to exhibit greater control over visual representations, especially those used information retrieval systems, has potential for both positive and negative impacts. On the positive side, it allows users greater control over visual information retrieval. This capability is particularly useful for visually-impaired users that can struggle to directly use and interpret images. However, introducing natural language control over visual representations has potentially negative impacts. We highlight two potential harms: (1) Accidental introduction of bias due to natural language conditioning. Large language models exhibit similar biases to humans (Guo et al., 2024). LLM representations used in information retrieval may adopt these biases. Furthermore, allowing users natural language control over visual representations may result in users accidentally introducing their own biases into visual retrieval. (2) Malicious control over information retrieval. Introducing natural language control over visual representations creates an entry-point to filter or modify what information is retrieved by downstream users. This is alarming in the internet age, where control over information access is a powerful tool for authoritarian control.
Appendix A Instruction Fine-tuning Prompt Design
@@
Appendix B CtrlBench Prompt Design
We use the following template in Gemini 2.5 Flash for each of the 1000 images:
"<Image> Your task is to identify and describe 5 distinct objects or elements in this image. For each object provide a short sentence description of it. Please be specific in your description. Then, for each description, provide a user prompt where that description caption would be a reasonable answer. Output your answer as JSON in the following template: {"desc_1": <DESCRIPTION 1>, "prompt_1": <PROMPT 1>, "desc_2": <DESCRIPTION 2>, "prompt_2": <PROMPT 2>, "desc_3": <DESCRIPTION 3>, "prompt_3": <PROMPT 3>, "desc_4": <DESCRIPTION 4>, "prompt_4": <PROMPT 4>, "desc_5": <DESCRIPTION 5>, "prompt_5": <PROMPT 5>}
Save on API costs, we generate all the candidates and queries for each in a single prompt.
Appendix C Architecture Ablation
LLM2Vec (BehnamGhader et al., 2024) introduced several architecture modifications for adapting LLMs into text embedding models. In the multimodal setting, we find that using casual vs. bidirectional attention is marginal for improving model accuracy (Figure˜7).
@@ Appendix C Architecture Ablation
On average, bidirectional attention slightly outperforms casual attention. Ablating over adapter rank, we find that different rank adapters tend to converge to the same accuracy, with higher rank adapters performing a fraction of a percentage better. Overall, better data and scaling (Appendix˜D) present a much more promising direction for improving embedding quality.
Appendix D Scaling Training
We find that batch size and number of samples seen during training are both important factors. By increasing our batch size by 44x, from $128$ queries and $1024$ candidates to our full pretraining run size ($512$ queries and $4096$ candidates), our validation accuracy on an $800$ sample validation batch using in-batch negatives increases from $92.5\%$ to $95.0\%$. We control for total samples seen by training the smaller batch size model for 44x more steps. Interestingly, this effect is even more pronounced on OOD validation data. For example, on a similar batch of MSCOCO data, accuracy increases from $73.8\%$ to $84.0\%$. Therefore, techniques for scaling batch size under limited VRAM, such as GradCache Gao et al. (2021), are a promising direction to improve our pretrained model. Furthermore, we find that more steps (more samples seen) is also a straightforward way to increase the pretrained model’s performance. As shown by Figure˜8, doubling step count steadily decreases loss during our pretraining run.
Figure 8: More samples seen vs. loss.
Appendix E Comparison to VLM2Vec
With the MMEB benchmark, Jiang et al. (2024) also include a mutltimodal embedding model: VLM2Vec. We exclude VLM2Vec from Tables˜1, 2 and 5 as it has been trained on MSCOCO and several tasks from the MMEB benchmark. We provide the comparison of the two models here while specifying which tasks are out of distribution (OOD) for VLM2Vec. All tasks are OOD for our model.
@@ Appendix E Comparison to VLM2Vec
HatefulMemes Kiela et al. (2021) 67.1 52.1
VOC2007 Everingham et al. 88.6 81.4
SUN397 Xiao et al. (2010) 72.7 71.8
Place365 López-Cifuentes et al. (2020) 42.6 40.7
@@
ObjectNet Barbu et al. (2019) 29.5 67.7
Country-211 Radford et al. (2021) 13.0 18.5
@@
Table˜7 compares the performance of our model to VLM2Vec Jiang et al. (2024). Gray rows indicate tasks that VLM2Vec is trained on. On average, VLM2Vec performs better on tasks that are in its distribution, while our model outperforms on tasks where both models are OOD. Interestingly, our model performs better on ImageNet-1K even though it is not trained on it.
Appendix F Results on VLLMs without Contrastive Training
Can dense embedding be extracted from VLLM’s without additional contrastive training? We experiment with two common methods for extracting dense embeddings from last-layer hiddens: (1) Using the last token in the sequence (Jiang et al., 2024). (2) Averaging over the sequence dimension (BehnamGhader et al., 2024).
Pooling Method Direction Top-1 Top-5 Top-10
Mean Pooling i2t 6.30% 16.18% 23.70%
t2i 3.41% 9.25% 13.39%
Last Token i2t 0.06% 0.26% 0.36%
t2i 0.08% 0.30% 0.73%
Table 8: Comparison of pooling methods across retrieval directions and on MSCOCO.
On MSCOCO retrieval we find that mean pooling significantly outperforms using the last token hidden (Table˜8). However, neither method performs very strong. Both methods are significantly worse than a zero-shot CLIP model on MSCOCO (Table˜1).
Appendix G Inference Settings for CLIP
Resolution Image $\rightarrow$ Text Text $\rightarrow$ Image
224x224 64.94 48.02
336x336 63.28 47.42
448x448 61.98 45.97
Table 9: OpenClip-ViT-g-14 accuracy (top-1) on MSCOCO at different image resolutions. Unlike ABC, accuracy degrades when scaling token budgets beyond OpenClip-ViT-g-14’s native $224\times 224$ resolution.
We find that CLIP models perform best at their pretraining resolution. As shown in Table˜9 accuracy consistently degrades when scaling the token counts of CLIP models beyond their native resolution. It would be unfair to evaluate models at a setting that is both more compute intensive and less performant. Therefore, we evaluate CLIP models at their native resolution and token count across all experiments.
Appendix H CtrlBench Examples